Brittany saves the same amount of money each week. She had $250 eight weeks ago. After withdrawing $10, her current balance is $400. How much was she saving weekly?
Answer by checkley71(8403)

250+8D-10=400
8D=400-250+10
8D=160
D=160/8
D=$20 IS THE WEEKLEY AMOUNT BRITTANY DEPOSITS.